Determine the equation of the line perpendicular to 3y + 4x = 1 and passing through the point A (-1, 4)
1
Expert's answer
2020-10-25T18:35:32-0400

"\bold {Answer}"


"4y -3x = 19"


"\bold {Solution}"

The the products of the gradients of perpendicular lines is "-1" , that is, "mm' = -1" where "m" is the gradient of one line and "m'" is the gradient of a line perpendicular to the first line.


The general form of a straight line equation is "y = mx + c" . Reducing "3y +4x = 1" to standard form gives:

"3y + 4x = 1"

"=> 3y = -4x + 1"


"=> \dfrac {3y}{3} = \dfrac {-4x}{3} + \dfrac {1}{3}"


"=> y = -\dfrac {4}{3}x + \dfrac {1}{3}"


Now, the gradient is "-\dfrac {4}{3}"


Since "mm' = -1"

"=> -\dfrac {4}{3}m' = -1"

"=> -4m' = -1 \u00d7 3"

"=> m' = \dfrac {-3}{-4}"


"= \dfrac {3}{4}"


Thus, the perpendicular line has gradient "\dfrac {3}{4}"


The equation of the perpendicular line through "\bold A" is given by:


"y - y_{1} = m'(x-x_{1})"

"=> y - 4 = \dfrac {3}{4} (x - (-1))"


"=> y = \dfrac {3}{4}(x+1)+4"


"=>4y = 3(x+1) + 16"


"=> 4y = 3x + 3+16"


"=> 4y -3x = 19"